# MicroMoE: Fine-grained Load Balancing for Mixture-of-Experts with Token Scheduling

### v1

@@ 9. Discussion
Scalability. Scaling MicroMoE to large clusters and increased expert counts introduces new challenges to the system. When scaling up, MicroMoE benefits from an expanded scheduling space, which enhances the load balancing capability of MicroEP. However, this larger scheduling space increases computational overhead in scheduling, potentially degrading the system efficiency. To compromise between load balancing capability and system efficiency, we can organize GPUs and experts into groups and perform scheduling at the group level, similar to prior works (SmartMoE; GShard; DeepSeek-V3).
@@
FSDP. We currently implement MicroMoE based on Megatron-LM’s DDP (Megatron-LM). Megatron-LM also supports Fully Sharded Data Parallel (FSDP) (FSDP), which resembles DeepSpeed’s ZeRO-3 (ZeRO), sharding model parameters and gradients for memory saving. Moreover, FSDP allows users to selectively enable sharding for model modules, providing a fine-grained control of ZeRO’s optimizations. We plan to integrate MicroMoE with FSDP in future work.
10. Conclusion
@@
We propose MicroEP, a novel expert parallelism strategy to achieve fine-grained load balancing in MoE. MicroEP dynamically balances GPU loads within every micro-batch through token scheduling. We primarily make two optimizations in MicroEP: First, we formulate the token scheduling process as a linear programming problem, which can be solved efficiently. Second, we theoretically analyze the relationship between expert placement and load balancing capacity and develop two placement strategies for different training scenarios. Finally, we propose MicroMoE, an efficient MoE training system based on MicroEP. Our experimental evaluation demonstrates that MicroMoE achieves significant performance improvements, with up to 47.6% end-to-end speedup compared to Megatron-LM, and almost consistently maintains optimal load balance across GPUs.
This work does not raise any ethical issues.
Appendix A Optimizations to MicroEP
A.1. Communication-Aware Scheduling
@@
Many studies demonstrate that the all-to-all communication also takes a significant amount of time in MoE layers (Tutel; FasterMoE; Lina). However, the optimization problem 1 only considers the computation time. Therefore, we can further consider the communication time during scheduling.
@@
In many existing frameworks (Megatron-LM; DeepSpeed), the time of an MoE layer is mainly determined by the communication time + the computation time (unless using some communication-computation overlapping techniques, such as Comet (Comet) and DualPipe (DeepSeek-V3)). Therefore, we introduce an additional objective to minimize the maximum communication volume on a GPU. According to Algorithm 1, Line 6, the local data volume of GPU $g$ can be formulated as $local_{g}=\sum\limits_{e\in E:g\in G_{EDP}^{e}}\min(x_{e}^{g},input_{e}^{g})$. For each GPU, we consider its all-to-all communication volume as the greater value between its send volume and receive volume. The send/receive volume of GPU $g$ can be calculated from $x_{e}^{g},input_{e}^{g}$, and $local_{g}$. Putting them together, we derive the new optimization problem:
(4) minimize $comp+\alpha\cdot comm,$
subject to $comp=\max\limits_{g\in G_{MicroEP}}\left\{\sum\limits_{e\in E:g\in G_{EDP}^{e}}x_{e}^{g}\right\},$
$comm=\max\limits_{g\in G_{MicroEP}}\left\{\max(send_{g},recv_{g})\right\},$
@@
A.2. Pipelining MicroEP
(a) MicroEP without pipelining.
(b) MicroEP with pipelining.
Figure 12. Pipelining MicroEP.
Pipelining offers an additional approach to hide the scheduling overhead of MicroEP, different from the overlapping technique in §5.4. Pipelining is particularly useful in scenarios where it is impossible to fully hide the scheduling time through overlapping. For example, some high-performance frameworks, such as DeepEP (DeepSeek-V3), have no intermediate operations to overlap between the gate network and all-to-all communication; when dealing with large-scale deployments with numerous GPUs and experts, the scheduling can incur substantial overhead (as illustrated in §7.5). With pipelining, we can overlap the scheduling of some tokens with the all-to-all communication of other tokens. This approach provides more flexibility in latency hiding.
We find many design choices to perform pipelining for MicroEP and illustrate one of them as follows. We observe that MicroEP can achieve perfect load balance in most circumstances (as shown in §7.3). Our intuition is that even if the expert load becomes more imbalanced, MicroEP can still achieve perfect balance. Inspired by this observation, we can split the tokens into two parts: We apply EP to the former part [^4] Since we have already changed expert placement, this EP is somehow different from typical EP and more like FlexMoE., and apply MicroEP to the latter part. Moreover, the optimization problem 1 should consider the computational workloads in both EP and MicroEP. In this way, we can overlap the scheduling of the latter part with the all-to-all communication of the former part, as shown in Figure 12.
Overhead. Pipelining MicroEP introduces some additional system overhead. For example, splitting one all-to-all communication operation into two increases synchronization time and GPU kernel launching time. Therefore, we recommend using MicroEP with pipelining in scenarios with substantial scheduling time but minimal system overhead.
@@ Appendix B Expert Placement using Cayley Graphs
B.1. Cayley Graphs
In many practical applications, we can construct near-optimal symmetric expert placements in MicroEP using Cayley graphs. The inherent symmetry of Cayley graphs makes them well-suited for constructing optimal expert placements.
@@ B.1. Cayley Graphs
We assume the MicroEP parameter $d=2$, so the hypergraph is a conventional graph. We observe that in practical applications, the quantities of GPUs and experts are usually powers of two. Let the number of GPUs be $2^{p}$, and the number of experts per GPU be $2^{q}$. Consequently, the number of vertices is $2^{p}$, the degree of each vertex is $2^{q}$, and the number of edges is $2^{p+q-1}$. In practice, we heuristicly construct many Cayley graphs for different $(p,q)$s.
@@
Figure 13. Examples of Cayley graphs.
B.2. Examples of Cayley Graphs
@@
B.3. Synchronization Consistency
Different EDP groups across experts can lead to a consistency issue during parameter and gradient synchronization. Specifically, the synchronization for different experts occurs in different EDP groups, which may incur deadlocks. To prevent deadlocks, we add a consistency restriction in expert placement: All replicas of an expert must have identical local expert indices. For example, in Figure 3(c), the replicas of expert 2 are the first local replicas of both GPU 1 and GPU 2; the colors of edges in Figure 13 also indicate the local expert indices. Since DDP executes parameter synchronization following the order of local parameters (and gradient synchronization following the reverse order) (DDP), deadlocks are effectively avoided.
@@ Appendix C Supplementary Experiments
C.1. Detailed Experimental Settings
@@
Activation recomputation is a technique to reduce memory footprint by avoiding recording activations in the forward pass and recomputing them in the backward pass (recomputation). Furthermore, selective activation recomputation enables recomputing a subset of model modules to perform a fine-grained trade-off between computation efficiency and memory (Megatron-SP). We enable selective recomputation in Megatron-LM to recompute only the MoE FFN, avoiding the out-of-memory (OOM) issue while maintaining relatively high throughput. Since DeepSpeed currently does not support selective recomputation, we recompute the whole layer in DeepSpeed. Furthermore, we find that we can adjust the granularity of selective recomputation at runtime. When the expert loads are highly imbalanced, we can recompute the whole MoE layer for better memory efficiency, avoiding OOM. Otherwise, we can recompute only the MoE FFN for better computation efficiency. For fair comparison, we do not adjust the recomputation granularity during evaluation.
Table 2. List of Model Hyperparameters
@@
C.2. Evaluation of DeepEP
Figure 14. Dispatch time of MicroEP and EP with DeepEP and NCCL, varying number of GPUs.
We evaluate the dispatch time of MicroEP and vanilla EP with DeepEP (DeepSeek-V3), a high-performance all-to-all communication backend. Megatron-LM (Megatron-LM) currently supports both NCCL (NCCL) (by default) and DeepEP for all-to-all communication. We additionally implement MicroEP with DeepEP.
Before analyzing the experimental results, two important experimental considerations should be noted: (1) Our testbed consists of 8 GPUs and only 2 NICs per node, resulting in limited inter-node network bandwidth. Therefore, we avoid employing EP or MicroEP across multiple nodes in §7. However, in this section, we focus on the performance of different communication backends, so we expand the communication group into multiple nodes. Consequently, the all-to-all time for inter-node communication is significantly longer than the intra-node communication. (2) In §7, we focus on the system performance, so we compare MicroEP using 8 GPUs per group with EP using 4 GPUs per group ($d=2$). However, in this section, we focus on the communication performance, so we compare MicroEP and EP using the same group size.
Figure 14 shows the dispatch time comparison between MicroEP and EP using both DeepEP and NCCL, varying number of GPUs. We use the same setting as in §7.4, except for the all-to-all group size. DeepEP exhibits better performance than NCCL due to its high-performance all-to-all implementation. When using NCCL, MicroEP requires less time than EP thanks to the locality-aware routing in §5.2. However, when using DeepEP, MicroEP requires more time than EP due to the data format incompatibilities between DeepEP and Megatron-LM. Consequently, Megatron-LM needs to pre-process the data for DeepEP, while MicroEP incurs a higher pre-processing overhead than EP.
C.3. Evaluation of Communication-Aware Scheduling
Figure 15. Execution time breakdown of an MoE layer with MicroEP, varying the levels of communication-aware scheduling.
@@ C.3. Evaluation of Communication-Aware Scheduling
C.4. Evaluation of MicroEP with Pipelining
Figure 16. Dispatch time breakdown with pipelining, vary the ratios of data in MicroEP (1.0 indicates no pipelining).
We evaluate the performance of MicroEP with pipelining in Appendix A.2. We enable the communication-aware scheduling and DeepEP. We use 8 GPUs and 128 experts. Other parameters are the same as Appendix C.3.
We compare the dispatch time with varying ratios of data in MicroEP. Figure 16 demonstrates that pipelining can reduce the dispatch time by overlapping MicroEP preparation with EP all-to-all communication. However, the dispatch time increases as the MicroEP ratio increases. This is because the EP all-to-all time decreases and becomes insufficient to fully hide the MicroEP scheduling time.

### v2

@@ 9 Discussion
Scalability. Scaling FineMoE to large clusters and increased expert counts introduces new challenges to the system. When scaling up, FineMoE benefits from an expanded scheduling space, which enhances the load balancing capability of FineEP. However, this larger scheduling space increases computational overhead in scheduling. To compromise between load balancing capability and system efficiency, we can organize GPUs and experts into groups and perform scheduling at the group level, similar to prior works [SmartMoE, GShard, DeepSeek-V3].
@@
FSDP. We currently implement FineMoE based on Megatron-LM’s DDP [Megatron-LM]. Megatron-LM also supports Fully Sharded Data Parallel (FSDP) [FSDP], which resembles DeepSpeed’s ZeRO-3 [ZeRO], sharding model parameters and gradients for memory saving. We plan to integrate FineMoE with FSDP in future work.
10 Conclusion
@@
We propose FineEP, a novel expert parallelism strategy to achieve fine-grained load balancing in MoE. FineEP dynamically balances GPU loads within every micro-batch through token scheduling. We primarily make two optimizations in FineEP: First, we formulate the token scheduling process as a linear programming problem, which can be solved efficiently. Second, we theoretically analyze the relationship between expert placement and load balancing capacity and develop two placement strategies for different training scenarios. Finally, we propose FineMoE, an efficient MoE training system based on FineEP. Our experimental evaluation demonstrates that FineMoE achieves significant performance improvements, with up to 47.6% end-to-end speedup compared to Megatron-LM, and almost consistently maintains optimal load balance across GPUs.
Appendix A Optimizations to FineEP
A.1 Communication-Aware Scheduling
@@
Many studies demonstrate that the all-to-all communication also takes a significant amount of time in MoE layers [Tutel, FasterMoE, Lina]. However, the optimization problem 1 only considers the computation time. Therefore, we can further consider the communication time during scheduling.
@@
In many existing frameworks [Megatron-LM, DeepSpeed], the time of an MoE layer is mainly determined by the communication time + the computation time (unless using some communication-computation overlapping techniques, such as Comet [Comet] and DualPipe [DeepSeek-V3]). Therefore, we introduce an additional objective to minimize the maximum communication volume on a GPU. According to Algorithm 1, Line 6, the local data volume of GPU $g$ can be formulated as $local_{g}=\sum\limits_{e\in E:g\in G_{EDP}^{e}}\min(x_{e}^{g},input_{e}^{g})$. For each GPU, we consider its all-to-all communication volume as the greater value between its send volume and receive volume. The send/receive volume of GPU $g$ can be calculated from $x_{e}^{g},input_{e}^{g}$, and $local_{g}$. Putting them together, we derive the new optimization problem:
minimize $comp+\alpha\cdot comm,$ (4)
subject to $comp=\max\limits_{g\in G_{FineEP}}\left\{\sum\limits_{e\in E:g\in G_{EDP}^{e}}x_{e}^{g}\right\},$
$comm=\max\limits_{g\in G_{FineEP}}\left\{\max(send_{g},recv_{g})\right\},$
@@
A.2 Pipelining FineEP
(a) FineEP without pipelining.
(b) FineEP with pipelining.
Figure 14: Pipelining FineEP.
Pipelining offers an additional approach to hide the scheduling overhead of FineEP, different from the overlapping technique in §5.4. Pipelining is particularly useful in scenarios where it is impossible to fully hide the scheduling time through overlapping. For example, some high-performance frameworks, such as DeepEP [DeepSeek-V3], have no intermediate operations to overlap between the gate network and all-to-all communication; when dealing with large-scale deployments with numerous GPUs and experts, the scheduling can incur substantial overhead (as illustrated in §7.6). With pipelining, we can overlap the scheduling of some tokens with the all-to-all communication of other tokens. This approach provides more flexibility in latency hiding.
We find many design choices to perform pipelining for FineEP and illustrate one of them as follows. We observe that FineEP can achieve perfect load balance in most circumstances (as shown in §7.3). Our intuition is that even if the expert load becomes more imbalanced, FineEP can still achieve perfect balance. Inspired by this observation, we can split the tokens into two parts: We apply EP to the former part [^5] Since we have already changed expert placement, this EP is somehow different from typical EP and more like FlexMoE., and apply FineEP to the latter part. Moreover, the optimization problem 1 should consider the computational workloads in both EP and FineEP. In this way, we can overlap the scheduling of the latter part with the all-to-all communication of the former part, as shown in Figure 14.
Overhead. Pipelining FineEP introduces some additional system overhead. For example, splitting one all-to-all communication operation into two increases synchronization time and GPU kernel launching time. Therefore, we recommend using FineEP with pipelining in scenarios with substantial scheduling time but minimal system overhead.
@@ Appendix B Expert Placement using Cayley Graphs
B.1 Cayley Graphs
In many practical applications, we can construct near-optimal symmetric expert placements in FineEP using Cayley graphs. The inherent symmetry of Cayley graphs makes them well-suited for constructing optimal expert placements.
@@ B.1 Cayley Graphs
We assume the FineEP parameter $d=2$, so the hypergraph is a conventional graph. We observe that in practical applications, the quantities of GPUs and experts are usually powers of two. Let the number of GPUs be $2^{p}$, and the number of experts per GPU be $2^{q}$. Consequently, the number of vertices is $2^{p}$, the degree of each vertex is $2^{q}$, and the number of edges is $2^{p+q-1}$. In practice, we heuristicly construct many Cayley graphs for different $(p,q)$s.
@@
Figure 15: Examples of Cayley graphs.
B.2 Examples of Cayley Graphs
@@
B.3 Synchronization Consistency
Different EDP groups across experts can lead to a consistency issue during parameter and gradient synchronization. Specifically, the synchronization for different experts occurs in different EDP groups, which may incur deadlocks. To prevent deadlocks, we add a consistency restriction in expert placement: All replicas of an expert must have identical local expert indices. For example, in Figure 3(c), the replicas of expert 2 are the first local replicas of both GPU 1 and GPU 2; the colors of edges in Figure 15 also indicate the local expert indices. Since DDP executes parameter synchronization following the order of local parameters (and gradient synchronization following the reverse order) [DDP], deadlocks are effectively avoided.
@@ Appendix C Supplementary Experiments
C.1 Detailed Experimental Settings
@@
Activation recomputation is a technique to reduce memory footprint by avoiding recording activations in the forward pass and recomputing them in the backward pass [recomputation]. Furthermore, selective activation recomputation enables recomputing a subset of model modules to perform a fine-grained trade-off between computation efficiency and memory [Megatron-SP]. We enable selective recomputation in Megatron-LM to recompute only the MoE FFN, avoiding the out-of-memory (OOM) issue while maintaining relatively high throughput. Since DeepSpeed currently does not support selective recomputation, we recompute the whole layer in DeepSpeed. Furthermore, we find that we can adjust the granularity of selective recomputation at runtime. When the expert loads are highly imbalanced, we can recompute the whole MoE layer for better memory efficiency, avoiding OOM. Otherwise, we can recompute only the MoE FFN for better computation efficiency. For fair comparison, we do not adjust the recomputation granularity during evaluation.
Table 2: List of Model Hyperparameters
@@
C.2 Evaluation of DeepEP
Figure 16: Dispatch time of FineEP and EP with DeepEP and NCCL, varying number of GPUs.
We evaluate the dispatch time of FineEP and vanilla EP with DeepEP [DeepSeek-V3], a high-performance all-to-all communication backend. Megatron-LM [Megatron-LM] currently supports both NCCL [NCCL] (by default) and DeepEP for all-to-all communication. We additionally implement FineEP with DeepEP.
Before analyzing the experimental results, two important experimental considerations should be noted: (1) Our testbed consists of 8 GPUs and only 2 NICs per node, resulting in limited inter-node network bandwidth. Therefore, we avoid employing EP or FineEP across multiple nodes in §7. However, in this section, we focus on the performance of different communication backends, so we expand the communication group into multiple nodes. Consequently, the all-to-all time for inter-node communication is significantly longer than the intra-node communication. (2) In §7, we focus on the system performance, so we compare FineEP using 8 GPUs per group with EP using 4 GPUs per group ($d=2$). However, in this section, we focus on the communication performance, so we compare FineEP and EP using the same group size.
Figure 16 shows the dispatch time comparison between FineEP and EP using both DeepEP and NCCL, varying number of GPUs. We use the same setting as in §7.4, except for the all-to-all group size. DeepEP exhibits better performance than NCCL due to its high-performance all-to-all implementation. When using NCCL, FineEP requires less time than EP thanks to the locality-aware routing in §5.2. However, when using DeepEP, FineEP requires more time than EP due to the data format incompatibilities between DeepEP and Megatron-LM. Consequently, Megatron-LM needs to pre-process the data for DeepEP, while FineEP incurs a higher pre-processing overhead than EP.
C.3 Evaluation of Communication-Aware Scheduling
Figure 17: Execution time breakdown of an MoE layer with FineEP, varying the levels of communication-aware scheduling.
@@ C.3 Evaluation of Communication-Aware Scheduling
C.4 Evaluation of FineEP with Pipelining
Figure 18: Dispatch time breakdown with pipelining, vary the ratios of data in FineEP (1.0 indicates no pipelining).
We evaluate the performance of FineEP with pipelining in Appendix A.2. We enable the communication-aware scheduling and DeepEP. We use 8 GPUs and 128 experts. Other parameters are the same as Appendix C.3.
We compare the dispatch time with varying ratios of data in FineEP. Figure 18 demonstrates that pipelining can reduce the dispatch time by overlapping FineEP preparation with EP all-to-all communication. However, the dispatch time increases as the FineEP ratio increases. This is because the EP all-to-all time decreases and becomes insufficient to fully hide the FineEP scheduling time.